OpenStudy (anonymous):

HELP! i dont know how to do this! You attend an amusement park with your family. Your parents buy you an all-ride pass for $20, shown as f(x). Instead of getting a pass, your parents decide to pay $4 for each ride they take, shown as g(x). What function shows the correct combination of these two functions to represent the total cost to them of attending the amusement park that day, shown as h(x)?

OpenStudy (anonymous):

do you have options to choose from?

OpenStudy (anonymous):

f(x) = 20x, g(x) = 4, h(x) = 20x + 4 f(x) = 20, g(x) = 4, h(x) = 4 + 20 f(x) = 20, g(x) = 4x, h(x) = 4x + 20 f(x) = 20x, g(x) = 4x, h(x) = 20x + 4x

OpenStudy (anonymous):

well f(x) is 20x and g(x) is 4x, to get h(x) you would add 20x + 4x

OpenStudy (anonymous):

so 24x

OpenStudy (anonymous):

it only asks for the function so it would be the last one

OpenStudy (nurali):

answer is the third one. see f(x)= 20, 20 is the cost of child's cost. g(x)=4x, x is the number of ticket they paid for. then h(x)=4x+20, the total cost

OpenStudy (anonymous):

thanks so much!!